When conducting a T-test, the difference between two groups is 1 and 0.337. What does the results say? Is 1 a significant difference and what does it mean?
Answer #1

Ans:

Given that

point estima for difference=sample means difference=1

standard error for difference=0.337

Test statistic:

t=1/0.337=2.967

say,df=18

(there are 10 subjects in each group,so df=10+10-2=18)

p-value=tdist(2.967,18,2)=0.0083

Assume significance level=0.01

As,p-value<0.01,we reject null hypothesis.

There is sufficient evidence to conclude that there is d=significant difference between means of two groups.